OpenStudy (anonymous):

alisha is making punch for a graduation party. The recipe requires twice as much orange juice as cranberry juice and 8 times as much ginger ale as cranberry juice. If she plans to make 165 oz of punch, how much orange juice does she need?

OpenStudy (asnaseer):

let 'o'=amount of orange juice, 'g'=amount of ginger ale and 'c'=amount of cranberry juice. we can then write: o = 2c g = 8c and the entire punch contains all three, so we can also write: o + g + c = 165 therefore: 2c + 8c + c = 165 11c = 165 c =15 oz therefore o=2c=2*15=30 oz